OpenStudy (anonymous):

Given the function f(x) = 5x - 9, find f(-3)

OpenStudy (anonymous):

put instead of x -3 f(-3)=5(-3)-9=...

OpenStudy (anonymous):

and

OpenStudy (anonymous):

you do the calculation

OpenStudy (anonymous):

wt bout f

OpenStudy (anonymous):

f(-3)=-15-9=-24